OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -4, 12, -28, 60, … A. -4, 12, -28, -60, … B. -4, -12, -28, -60, … C. -4, 12, -18, 54, … D. -4, 12, -18, -54, …

OpenStudy (anonymous):

Please Help Me!! ASAP

OpenStudy (whpalmer4):

So according to your formula, \[f(2) = -3*f(2-1) - 2*f(2-2)\]and\[f(0)=-4, \,f(1) = 12\] Can you find the value of \(f(2)\)?

OpenStudy (anonymous):

No I am very confused by this can you please help

OpenStudy (whpalmer4):

Hint: \[f(2-1) = f(1)\]and\[f(2-2) = f(0)\]

OpenStudy (anonymous):

I still have no idea my teacher barely went over this

OpenStudy (whpalmer4):

Look, it's simple substitution. Everywhere you see \(f(1)\), put (12). Everywhere you see \(f(0)\), put (-4).

OpenStudy (anonymous):

Then what would i do with this f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (anonymous):

WAIT WOULD THE ANSWER be C because it cant be a or be because it is 12 not -12

OpenStudy (anonymous):

Please HELP ME

OpenStudy (mathmale):

given that f(0) = -4 and f(1) = 12, and that the recursive formula is f(n) = -3•f(n -1) - 2•f(n - 2), let's find f(2). f(2) = -3*f(2-1) - 2f*(2-2) = -3*f(1) - 2*f(0) Would you now substitute the given values for f(0) and f(1) to calculate f(2).

OpenStudy (anonymous):

What i am still very confused so it would end up being -50

OpenStudy (anonymous):

Can you please help am i correct or can you just tell me the answer because i have been stuck on this for like half an hour

OpenStudy (whpalmer4):

Everywhere you see \(f(1)\) replace it with \((12)\). Everywhere you see \(f(0)\) replace it with \((-4)\). \[f(2) = -3*f(1) - 2*f(0)\] What do you get after you do those replacements?

OpenStudy (anonymous):

f(2) = 152?

OpenStudy (anonymous):

oh wait i did something wrong

OpenStudy (anonymous):

f(2) = -28 right?

OpenStudy (whpalmer4):

\[f(2) = -3*f(1) - 2*f(0)\]\[=-3*(12) - 2*(-4)\]\[=-36-(-8) \]\[= -28\]

OpenStudy (whpalmer4):

See, that's not so hard, is it? Now to find \(f(3)\) we do the same thing, except here we use \(f(2) = -28\) and \(f(1) = 12\) instead of \(f(1) = 12\) and \(f(0) = -4\)

OpenStudy (whpalmer4):

\[f(3) = -3*f(2)-2*f(1) =\]

OpenStudy (anonymous):

ok so the answer would be A

OpenStudy (whpalmer4):

let's just work through it before making any rash decisions about the answers, okay?

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

so \(f(3) =\)

OpenStudy (anonymous):

60

OpenStudy (whpalmer4):

it looks like you might have copied the answers wrong in the problem statement... Yes, that's right. What is \(f(4)=\) (yes, you don't need it for this problem, but I'd like to see that you can find it without our help!)

OpenStudy (anonymous):

-124

OpenStudy (whpalmer4):

very good! So you mentioned that you thought A was the right choice. Looking at what you've posted, you have A listed as \[\text{A}. -4, 12, -28, -60, … \] I assume that is really supposed to be \(60\) on the end?

OpenStudy (anonymous):

yes sorry i copied it wrong

OpenStudy (anonymous):

While Your Still here could you help me with another question

OpenStudy (anonymous):

The first four terms of a sequence are shown below. 9, 5, 1, -3 Which of the following functions best defines this sequence? f(1) = 9, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 9, f(n + 1) = f(n) + 4; for n ≥ 1 f(1) = 9, f(n + 1) = f(n) - 5; for n ≥ 1 f(1) = 9, f(n + 1) = f(n) + 5; for n ≥ 1